OpenStudy (anonymous):

x + 5y = -2 2x + y = 5 The point of intersection of the lines has a y-coordinate of _____. -1 1 3

OpenStudy (anonymous):

x+5y = -2 -10x+5y = 25 -9x = -27 x = 3 Then sub 3 back in to x + 5y = -2 5y = -5 y = -1

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

no problem do you understand what I did?